OpenStudy (anonymous):

Someone help me please. I'm trying so hard to understand this. A bank account has a beginning balance of $560.00. After 6 months, the balance in the account has increased to $572.60. What interest rate is being earned annually on this account? .045% 2.20% 2.25% 4.5%

OpenStudy (mathstudent55):

How much interest was earned in the 6 months? I'm not talking about the interest rate, I mean, how much money in dollars was earned in interest in the 6 months?

OpenStudy (anonymous):

$12.60?

OpenStudy (mathstudent55):

Yes, exactly. It's the difference between the amount of money before and after the 6 months passed. $572.60 - $560 = $12.60

OpenStudy (mathstudent55):

$12.60 in interest was earned in 6 months. At the same rate, how much would be earned in 1 year.

OpenStudy (anonymous):

Ohhh, I see now.

OpenStudy (anonymous):

But wait, I need to find the percentage...

OpenStudy (mathstudent55):

One step at a time. What is the total interest that would be earned in 1 year at this rate?

OpenStudy (mathstudent55):

Are you still there?

OpenStudy (anonymous):

I'm here.. I'm confused.

OpenStudy (mathstudent55):

No need for confusion. Let's make it clear. In 6 months, $12.60 in interest was earned. 1 year is 2 times 6 months, so in 1 year, two times that interest would be earned. What is 2 * $12.60 = ?

OpenStudy (anonymous):

25.2

OpenStudy (mathstudent55):

Correct. Now there is only one step left. We need to find out what percentage 25.20 is of 560.

OpenStudy (anonymous):

4.5%?

OpenStudy (mathstudent55):

We can do that with a proportion: 560 is to 100% as 25.20 is to x% \(\dfrac{560}{100} = \dfrac{25.20}{x}\)

OpenStudy (mathstudent55):

We cross multiply: \(560x = 25.20 \times 100\) \(560x = 2520\) \(x = \dfrac{2520}{560} \) \(x = 4.5\) Answer: 4.5%

OpenStudy (mathstudent55):

You are correct.

OpenStudy (mathstudent55):

The reason that our second step was to multiply $12.60 by 2 is that the $12.60 in interest if for 6 months. The question is asking what the annual interest rate is, so we needed to know how much interest would be earned in 1 year.

OpenStudy (mathstudent55):

Good job! Bye